OpenStudy (anonymous):

In which direction must the graph of f(x) = x2 be shifted to produce the graph of g(x) = (x + 7)2?

OpenStudy (inkyvoyd):

Now, consider the equation y=a(x-h)^2+k Recognize it? It's vertex form!

OpenStudy (anonymous):

left

OpenStudy (inkyvoyd):

Now, for f(x), (h,k) is (0,0), because f(x)=1(x-0)^2+0

OpenStudy (inkyvoyd):

g(x)=1(x+7)^2

OpenStudy (inkyvoyd):

the vertex of g(x) is (-7,0)

OpenStudy (inkyvoyd):

Because the vertex of g(x) is to the left of the vertex of f(x), we must shift f(x) left.

OpenStudy (anonymous):

ok,, i'll work this out on my sheet of paper so i can have a definate understanding,, it'll only take a couple of minutes !

OpenStudy (anonymous):

RIGHT AGAIN !!

OpenStudy (inkyvoyd):

Does it all make sense to you?

OpenStudy (anonymous):

yes it does....lol well now it does !